OpenStudy (anonymous):

Gary is using a combination of regular polygonal tiles for his kitchen floor. He begins the tessellation at a vertex by placing one square and two triangles. Which two polygons should he use to fill up the vacant space at the vertex and make it possible to tessellate the floor?

OpenStudy (anonymous):

a.) 2 hexagons because 360 – (120 + 60 +60) = 120 b.) 2 squares because 180 = 90 + 90 c.) 1 triangle and 1 hexagon because 180 = 60 + 120 d.) 1 square and 1 triangle because 360 – (90 + 60 + 60) = 90 + 60

OpenStudy (anonymous):

@mathstudent55 Can You Help Me With This Question? PLEASE :D

OpenStudy (anonymous):

@mathmate Can You Help Me PLEASE?

OpenStudy (mathmate):

I am looking at the question and not too sure what the two triangles would be like. There are many ways to put two triangles.

OpenStudy (goldphenoix):

Yeah, I'm not sure. It's not clear. Then again, I don't remember a lot about tessellation.

OpenStudy (anonymous):

Can It Be Two Hexagons? @GoldPhenoix @mathmate

OpenStudy (goldphenoix):

I am not sure, sorry. =/

OpenStudy (mathmate):

 This could be one way.